Circular No. 7693

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dy IN MCG +04-40-16 T. Puckett and M. Marcus, Mountain Town, GA, report the discovery

of an apparent supernova (mag 18.0) on an unfiltered CCD frame (limiting mag 19.4) taken with the Puckett Observatory 0.30-m automated patrol telescope on Aug. 27.12 UT. SN 2001dy is located at α = 17^{h}04^{m}59^{s}.42, δ = +23^{o}10′06′′.2 (equinox 2000.0), which is 5′′.9 west and 2′′.8 south of the center of MCG +04-40-16. The new object was confirmed on CCD frames taken on Aug. 28.07 by Puckett with a 0.60-m reflector. SN 2001dy is not present on images taken by Puckett in 1999 (Aug. 10 and 17) and 2000 (Mar. 31, Apr. 30, May 9, and June 2; limiting mag ∼ 20.5), neither is it present on Palomar Sky Survey images taken on 1992 Sept. 7, 1990 Apr. 28 (limiting mag ∼ 21.0), and 1955 Apr. 1 (limiting mag ∼ 19.7).

SUPERNOVA 2001dr IN NGC 4932 S. Benetti, G. Altavilla, E. Cappellaro, A. Pastorello, and M. Turatto,

Astronomical Observatory, Padua; and A. Della Valle, Astronomy Department, Padua, write: “A reduced CCD spectrum (range 380–770 nm, resolution 2.5 nm) of SN 2001dr (cf. IAUC 7684), obtained on Aug. 25.83 UT with the Asiago 1.82-m telescope (+ AFOSC), is that of a type-II supernova, ∼ 1 week after maximum light. The spectrum shows a broad Hα emission with a P-Cyg absorption having a minimum measured at 649.7 nm. Adopting the NASA/IPAC Extragalactic Database recession velocity for NGC 4932 of 7088 km/s, we find an expansion velocity of ∼ 9800 km/s.”

V857 CENTAURI J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and

Space Research Organization Netherlands (SRON), Utrecht; R. Mewe and J. Heise, SRON, write: “Analysis of archival data from the Wide Field Cameras on BeppoSAX has revealed an energetic x-ray flare from a position consistent with that of V857 Cen = Gliese 431, a dM4.5e star (V = 11.5) at a distance of 10.5 pc that was seen to flare once before in the optical (Doyle et al. 1990, A.Ap. Suppl. 86, 403). The x-ray flare occurred on May 28.739 UT, had a rise time of 1 to 3 min, and lasted at least 12 min. After a data gap from 12 to 88 min after the onset, the flare was over. The peak luminosity is 3 × 10^{31} erg s^{−1} cm^{−2} (2–28 keV), and the fluence at least 2 × 10^{34} ergs. The spectral shape is consistent with thermal bremsstrahlung at 11 ± 4 keV. This is the only flare detected from Gl 431 above 1 × 10^{31} erg s^{−1} cm^{−2} in 1 month of net exposure time over the past five years.”

2001 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green